OpenStudy (anonymous):

rewarded medal and fan A blueprint of a shopping complex shows the bottom edge of the roof to be 93 feet above the ground. If the roof rises to a point 171 feet above the ground over a horizontal distance of 6.5 yards, what is the slope of the roof?

OpenStudy (anonymous):

have you tried drawing a picture of what this might look like?

OpenStudy (anonymous):

no,,, i have no idea how to do this

OpenStudy (anonymous):

oh yes

OpenStudy (anonymous):

but how do I get the equation?

OpenStudy (anonymous):

ok so the equation for the slope is always this: \[slope = \frac{rise}{run} = \frac{ \Delta y}{\Delta x}\] does this look familiar?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

would it be 171ft-93ft/19.5ft?

OpenStudy (anonymous):

alright so we got to find the change in the height of the roof and divide it by the width of the roof...btw I just realized that I drew it wrong lol...(it should not have two parts but just one) AND YES but you need the parentheses so....(171-93)/19.5

OpenStudy (anonymous):

here's the corrected picture so I don't confuse you: 

OpenStudy (anonymous):

so would the answer be 8??

OpenStudy (anonymous):

haha close so... 171 -93 = 78

OpenStudy (anonymous):

78/19.5 is what?